# Programmer Visual Attention During Context-Aware Code Summarization

### v1

@@ 3.9 Participants
We recruited 10 programmers with Java experience to participate in our study. Each participant was required to have at leaszt 1 year of previous Java development experience. The participants had an average of 2.8 years of Java development experience and 5 years of general programming experience. Each participant was compensated at a flat rate of USD 60 per session, with an average duration of around 1 hour per session. 4 participants identified as female, while 6 identified as male. 6 participants identified as non-native but fluent English speakers, while 4 identified as native English speakers. We designed the study to be in-depth to answer RQ1, where each participant processes multiple methods and projects, so we did not recruit a high number of participants. In this study we focus on general findings instead of individual differences.
@@ 4 Methodology
In this section, we discuss the methodology for analysis of data after the eye-tracking study.
TABLE I: Mean and Median values for metrics when grouped by the position in which the method was seen.
Method Fixation Count Fixation Duration Regression Rate Lines Visited Methods Visited
Mean Median Mean Median Mean Median Mean Median Mean Median
1 331.61 259 444.39 289 56.65 56.93 50.03 38.5 9.04 5
2 268.54 249.5 480.37 300 53.78 53.48 46.92 45 9.71 6.5
3 302.89 221 441.05 281 55.07 56.94 45.45 38 9.11 7
4 290.46 250.5 484.24 302 56.21 59.21 40.58 34 9.35 8
5 238.77 216 491.90 310.5 52.89 52.39 40.58 36.5 10.08 7
6 202.81 156 458.15 283 51.1 49.67 36.03 32 8.48 6
7 222.44 193 478.40 301.5 51.52 54.78 35.08 31 6.96 7
8 214.69 176 491.89 308 49.85 51.97 40.81 30.5 8.19 6
@@ 4.1 Gaze Pattern Metrics
A fixation is marked by a steady gaze of the human eye at a particular location in the stimulus, for the duration of at least 100-200ms, required for the visual system to process information. It is the most common measure of human visual attention in eye-tracking literature [44]. For this study, we only compute statistical summaries of fixation count and duration.
@@ 4.2 Qualitative Annotation
When assigning annotations or “codes” in a qualitative manner, there are two popular approaches to reduce human biases. The first approach is an agreement-based approach [49], which relies on a high number of annotators (or graders for the purpose of this study) to reach a Kappa agreement threshold. However, that approach does not account for expertise, i.e., a large number of novices may override an expert. For our study, every grader would have to be 1) expert in Java documentation, and 2) familiar with the entirety of the project, as we expect our participants to explore the project while writing summaries. The second approach is a discussion-based approach, which relies on an iterative process of discussion for each disagreement. We chose this approach as recommended by related work for tasks with a low tolerance for imperfect grading [50, 51]. Both graders independently graded all summaries, followed by a discussion phase on each disagreement, until both graders agreed on one score. Note that the graders did not know the participant number or the position of the method for which the summary was generated to eliminate agreement bias.
@@ 4.3 Context Categorization
We use a SRCML [42] archive to identify types of context for each fixation recorded in iTrace-toolkit [38]. Additionally, we use the Call-Hierarchy Plugin for Eclipse for Java IDE 2023 to generate the call graph context.
@@ 5 Results
In this section, we report the results for our three RQs.
@@ 5.1 RQ1: Effect of the progress through tasks
TABLE II: Results of the Mann-Whitney U statistical tests on the distributions presented in Table 1
@@
TABLE III: Average quality ratings for summaries when grouped by the position of the methods.
@@
TABLE IV: Gaze metrics for summaries grouped by quality scores: low(<=3) and high(=5) for completeness and conciseness. Here U1, U2, and p-value are results of the Mann-Whitney U statistical test for each group.
Group Fixation Count Regression Rate Fixation Duration Lines Visited Methods Visited
@@
We found that participants need less information to complete their context-aware summarization task as they progress through the tasks. To analyze the change in gaze patterns during this process, we report the change in fixation duration, fixation count, regression rate, lines visited, and methods visited by the participants. We make two interesting observations.
First, we observe a pattern of decreasing fixation counts. In Table I we report the statistical summaries of the  count of fixations when grouped by position of the method, which denotes the position in which the participants saw and wrote summaries for the method. We observe a 35% decrease in the mean number of fixations the participants made on the last method when compared to the first method processed. However, with respect to fixation duration, we did not observe a significant change. This may be because average time spent on a fixation may be affected by the reading and comprehension speed of the participants.
@@
We performed the Mann-Whitney U [53] test, a non-parametric, non-paired statistical test on the distribution of the values, and report the results in Table II. We chose this test because our data does not meet the assumptions of a parametric test, i.e., it is subjective and is not known to follow Gaussian or Normal distributions. We compared the combined distribution of first two methods against the combined distribution of last two methods seen by the participants. We observed that for fixation counts the difference between the two distributions were statistically significant (p-value$<0.01$). We also observed statistically significant difference between the two distributions for regression rate (p-value$=0.02$)
However, with respect to number of lines visited and number of methods visited, we observe a p-value$>0.05$. Therefore we cannot reject the null hypothesis for these metrics. There is statistical likelihood that patterns such as decreasing mean number of lines visited may not be significant. These observations are interesting because they indicate that the participants continued to cover a large amount of the context to write these descriptions as they processed more methods.
Therefore, our interpretation of these observations is that the participants retain more information about the project from previous task, and thus needed less information to write summaries as they processed more tasks. Different studies have shown how programmers skim source code whenever possible, reading the minimum amount of information they need [54]. As programmers become more familiar with the task, they are more likely to intuitively know what information they need, and therefore read fewer details. Our observations support those findings as we do not see a significant decrease in lines and methods visited by the participants but fixations and regressions decrease. Therefore, participants still covered similar amount of code but seemed to get better at skimming, resulting in lower frequency of fixations and regression rate.
@@ 5.2 RQ2: Quality of Summaries
First, we observe no significant pattern associated with the position in which the summary was written. In Table III, we report the mean scores for the four quality metrics. We also report an over score, an average of those 4 metrics calculated individually for each summary. We also performed the Mann-Whitney U test between the distributions of the first two methods and the last two methods in a way consistent with RQ1, with an observed $p>=0.9$. This observation supports our hypothesis about RQ1 that a decrease in fixation counts and regression rate is likely not due to exhaustion, as neither code context coverage nor quality of summaries was affected by the position in which the methods were seen.
@@
[…]
We observe that the completeness of summaries is correlated to the code context coverage by the programmer. In Table IV we report the mean values of each of the five gaze metrics for each subset. We also report the U1, U2, and p-value for Mann-Whitney U statistical test. We observe that the summaries of high completeness correspond to a significantly ($p\textless{}0.001$) higher number of fixations, lines visited, and methods visited. Figure 3 shows this delta clearly through the blue bars. This suggests that the participants that covered a smaller amount of the code context, wrote summaries that were missing important information. Although we observe an increase in mean regression rate as well, the p-value for regression rate $p=0.051$, which is not statistically significant by conventional standards, but approaches significance. We do not observe a significant difference in fixation duration. This observation supports the observation in RQ1, and further suggests that fixation duration may be dependant on each participant’s reading and comprehension speed.
On the other hand, the conciseness of summaries is inversely correlated to the code context coverage by the programmer. From Table IV, we observe that the summaries of high conciseness correspond to a significantly ($p<0.015$) lower number of fixations, lines visited, and methods visited. Figure 3 shows this delta more clearly. This suggests that participants included unnecessary information at a higher rate when they covered an exceptionally larger part of the code context. As shown in Figure 3, we did not observe a significant difference in fixation duration or regression rate between the two groups.
Our interpretation of these observations is that there are diminishing returns in terms of increasing code context coverage and improvement in the quality of summaries. We found that gaze metrics for the summaries that achieved a perfect score for all four qualities ($Overall=5$), tend to fall between the thresholds for high completeness and conciseness. For example, in terms of lines visited, the summaries with a perfect score have a mean score of 41 lines, which is between the scores of 43.25 and 38.49 reported in Table IV.
@@ 5.3 RQ3: Attention to Context
We found that participants spent a vast majority of their time looking at methods and variable declarations in same class as the target method. Surprisingly, we found that participants spent a comparatively small amount of time looking at methods in the call graph, even though we designed the second and third sentences in the summary to contain information from the callees and callers respectively. In Figure 2 we report four graphs, each showing averages for total time spent fixating on tokens inside each type of context, normalized by total time spent looking at context for the summarization task.
In Figure 2 (a), we report the attention grouped by the project type. We found see similar trends for four of the five project, namely scrimage, openaudible, mltk, and freecol. For these projects, participants spent an average of 40-65% of their time looking at methods in the same class as the target method. Surprisingly, for the mallet project, on an average participants spent more time looking at the variable, declarations, and other data inside the class than methods and calls. One possible explanation for this outlier we investigated was that mallet relies more on class variables than the other projects. However, we did not find any significant difference between average number of class declarations and variables outside class methods for mallet, when compared against the other projects.
In Figure 2 (b), we analyze the difference in attention grouped by session, i.e., the order in which the project was processed by a participant. We found that on average as participants processed more sessions, they relied more on information inside the class, both class methods and class declarations. This is expected because as participants repeat the task, they form a consistent pattern of navigating the context. This supports the earlier observation that overall, the participants found the class of the target method to be most important part of the context.
@@
In Figure 2 (d), we analyze the difference in attention of each participant. We found that most of the participants relied on class methods and the class declaration for contextual information. We found that only participant 5 and 6 are outliers, and spent a significant amount of time fixating on context in other files, and these are areas that do not fall under the methods in the call hierarchy. This suggests that observations from (a),(b), and (c) are consistent with a majority of the participants and not skewed heavily by an extreme minority.
@@ 6 Example
mallet>types>Multinomial.java>ensureCapacity [line 290]
     protected void ensureCapacity (int index)      {    if (index > size)  size = index;  if (counts.length <= index) {  @Var  int newLength = ((counts.length < minCapacity)           ? minCapacity           : counts.length);  while (newLength <= index)  newLength *= 2;  double[] newCounts = new double[newLength];  System.arraycopy (counts, 0, newCounts, 0,                          counts.length);  this.counts = newCounts;  }  }
participant 3
@@
participant 7
     This method is for ensuring that the Estimator class has  enough capacity to execute given the index that the method  is passed through. If index is great than the size of the  class, then the size is set to the index, if the length of  counts is less than the index, then it iterates through  counts to create new counts that is of length of the index.  This method is used by the project for the multinomial  class to make an estimate.
participant 10
@@
Figure 4: Example with source code of a method in project mallet, accompanied by summaries written by participant 3,7, and 10.
In Fig 4 we show an example containing target method and three corresponding summaries written by participant 3,7, and 10. These may explain some of our observations from Section 5. Notice, the summaries written by Participant 3 and 10 are similar, but the summary written by 10 is slightly clearer. On the other hand, the summaries written by Participant 7 for the same method are inaccurate and hard to comprehend and received one of the lowest overall scores by our graders (Overall = 2).
We found that the key to writing the last sentence for both participant 3 and participant 10’s summary is to track the class variable minCapacity. Instances like this might explain why mallet shows a higher attention on class declarations than class methods as observed in Section 5.3.
We also notice that participant 7’s summary is an example of summaries where excessive code coverage leads to verbose and unclear summary. Note this is just one example we found that fits the observations from Section 5.2, we do not claim that this example is representative of a trend or make any recommendations based on this example.
@@ 7 Threats to Validity
Another threat lies with our pool of participants. Our participants are graduate students, and their experiences may not translate directly to those of active industry professionals. To mitigate this threat, we screened participants with previous (and recent) Java development experience. However, industry professionals may have specific mental models for their industry and may utilize visual attention strategies that our participants do not represent. Another threat with our pool of participants is that half of them wear corrective glasses. In specific circumstances, corrective glasses can lead to false fixation readings. To try to mitigate this threat, we recommended that participants complete the task without glasses or use contact lenses if they are able.
@@
Another threat lies with the study conditions. The room in which we conducted the study has a window. As the study was conducted over several months, the lighting conditions may have changed between sessions. Newer generation eye-tracker claim to work in all lighting conditions but can cause inconsistencies in rare situations. These external threats exist for all eye-tracking studies.

### v2

@@ 3.9 Participants
We recruited 10 programmers with Java experience to participate in our study. We recruited participants by email list, and we collected informed consent which clearly indicated they could leave at any time, and were compensated 60 USD per session. Each participant was required to have at least 1 year of previous Java development experience. The participants had an average of 2.8 years of Java development experience and 5 years of general programming experience. Each participant was compensated at a flat rate of USD 60 per session, with an average duration of around 1 hour per session. 4 participants identified as female, and 6 as male. 6 participants identified as non-native but fluent English speakers, and 4 as native English speakers. We designed the study to be in-depth to answer RQ1, where each participant processes multiple methods and projects, so we did not recruit a high number of participants. In this study we focus on general findings instead of individual differences.
@@ 4 Methodology
In this section, we discuss the methodology for analysis of data after the eye tracking study.
@@ 4.1 Gaze Pattern Metrics
A fixation is marked by a steady gaze of the human eye at a particular location in the stimulus, for the duration of at least 100-200ms, required for the visual system to process information. It is the most common measure of human visual attention in eye tracking literature [49]. For this study, we only compute statistical summaries of fixation count and duration.
@@ 4.2 Qualitative Annotation
When assigning annotations or “codes” in a qualitative manner, there are two popular approaches to reduce human biases. The first approach is an agreement-based approach [54], which relies on a high number of annotators (or graders for the purpose of this study) to reach a Kappa agreement threshold. However, that approach does not account for expertise, i.e., a large number of novices may override an expert. For our study, every grader would have to be 1) expert in Java documentation, and 2) familiar with the entirety of the project, as we expect our participants to explore the project while writing summaries. The second approach is a discussion-based approach, which relies on an iterative process of discussion for each disagreement. We chose this approach as recommended by related work for tasks with a low tolerance for imperfect grading [55, 56]. Both graders independently graded all summaries, followed by a discussion phase on each disagreement, until both graders agreed on one score. To avoid agreement bias, the graders did not know the participant number or the order in which the summaries were shown to the participants.
@@ 4.3 Context Categorization
We use srcML [47] to automatically identify the type of source code element – such as conditional or method call – in each project. iTrace [43] calculates the file, column, and row data from the eye tracking data, which we use to identify the source code element the participant was looking at. From this we use srcML to identify the area of context in which the source code element is at each point.
@@ 5 Results
We report the results for our three RQs in this section. For perspective, key statistics about the dataset about the dataset include participants, gazes, and fixations per task:
Participants 10 Fixations (Mean) 263.5
Hours/Participant 6 Fixations (Min) 12
Total Fixations 95924 Fixations (Max) 1726
Total Gazes 143570
TABLE II: Mean and Median values for metrics when grouped by the position in which the method was seen.
Method Fixation Count Avg. Fixation Duration Regression Rate Lines Visited Methods Visited
Mean Median Mean Median Mean Median Mean Median Mean Median
1 331.61 259 444.39 289 56.65 56.93 50.03 38.5 9.04 5
2 268.54 249.5 480.37 300 53.78 53.48 46.92 45 9.71 6.5
3 302.89 221 441.05 281 55.07 56.94 45.45 38 9.11 7
4 290.46 250.5 484.24 302 56.21 59.21 40.58 34 9.35 8
5 238.77 216 491.90 310.5 52.89 52.39 40.58 36.5 10.08 7
6 202.81 156 458.15 283 51.1 49.67 36.03 32 8.48 6
7 222.44 193 478.40 301.5 51.52 54.78 35.08 31 6.96 7
8 214.69 176 491.89 308 49.85 51.97 40.81 30.5 8.19 6
@@ 5.1 RQ1: Effect of the progress through tasks
We divide our discussion of the answer to RQ1 into the following topics: Overall Information, Key Observations, Statistical Tests, and Interpretation.
TABLE III: Results of the Mann-Whitney U statistical tests on the distributions presented in Table  II
@@
TABLE IV: Average quality ratings for summaries when grouped by the position of the methods.
@@
TABLE V: Gaze metrics for summaries grouped by quality scores: low(<=3) and high(=5) for completeness and conciseness. Here U1, U2, and p-value are results of the Mann-Whitney U statistical test for each group.
Group Fixation Count Regression Rate Avg. fixation duration Lines Visited Methods Visited
@@
Overall Information We found that participants need less information to complete their context-aware summarization task as they progress through the tasks. To analyze the change in gaze patterns during this process, we report the change in average fixation duration, fixation count, regression rate, lines visited, and methods visited by the participants. We make two interesting observations. We found that participants need less information to complete their context-aware summarization task as they progress through the tasks. To analyze the change in gaze patterns during this process, we report the change in average fixation duration, fixation count, regression rate, lines visited, and methods visited by the participants.
Key Observations First, we observe a pattern of decreasing fixation counts. In Table II, we report the statistical summaries of the  count of fixations when grouped by position of the method, which denotes the position in which the participants saw and wrote summaries for the method. We observe a 35% decrease in the mean number of fixations the participants made on the last method when compared to the first method processed. However, with respect to average fixation duration, we did not observe a significant change. This may be because average time spent on a fixation may be affected by the participants’ reading speed.
@@
Statistical Tests We performed the Mann-Whitney U [58] test, a non-parametric, non-paired statistical test on the distribution of the values, and report the results in Table III. We chose this test because our data does not meet the assumptions of a parametric test, i.e., it is subjective and is not known to follow Gaussian or Normal distributions. We compared the combined distribution of the first two methods against the combined distribution of the last two methods seen by the participants. We observed that for fixation counts the difference between the two distributions were statistically significant ($p<0.01$). We also observed statistically significant difference between the two distributions for regression rate ($p=0.02$)
However, with respect to number of lines visited and number of methods visited, we observe a p-value $>0.05$. Therefore, we cannot reject the null hypothesis for these metrics. There is statistical likelihood that patterns such as decreasing mean number of lines visited may not be significant. These observations are interesting because they indicate that the participants continued to cover a large amount of the context to write these descriptions as they processed more methods.
Interpretation Our interpretation of these observations is that the participants retain more information about the project from previous task, and thus needed less information to write summaries as they processed more tasks. Different studies have shown how programmers skim source code whenever possible, reading the minimum amount of information they need [59]. As programmers become more familiar with the task, they are more likely to intuitively know what information they need, and therefore read fewer details. Our observations support those findings as we do not see a significant decrease in lines and methods visited by the participants but fixations and regressions decrease. Therefore, participants still covered similar amount of code but seemed to get better at skimming, resulting in lower frequency of fixations and regression rate.
@@ 5.2 RQ2: Quality of Summaries
First, we observe no significant pattern associated with the position in which the summary was written. In Table IV, we report the mean scores for the four quality metrics. We also report an over score, an average of those 4 metrics calculated individually for each summary. We also performed the Mann-Whitney U test between the distributions of the first two methods and the last two methods in a way consistent with RQ1, with an observed $p\geq 0.9$. This observation supports our hypothesis about RQ1 that a decrease in fixation counts and regression rate is likely not due to exhaustion, as neither code context coverage nor quality of summaries was affected by the position in which the methods were seen.
[…]
@@
We observe that the completeness of summaries is correlated to the code context coverage by the programmer. In Table V, we report the mean values of each of the five gaze metrics for each subset. We also report the U1, U2, and p-value for Mann-Whitney U statistical test. We observe that the summaries of high completeness correspond to a significantly ($p<0.001$) higher number of fixations, lines visited, and methods visited. Figure 3 shows this delta clearly through the blue bars. This suggests that the participants that covered a smaller amount of the code context, wrote summaries that were missing important information. Although we observe an increase in mean regression rate as well, the p-value for regression rate $p=0.051$, which is not statistically significant by conventional standards, but approaches significance. We do not observe a significant difference in average fixation duration. This observation supports the observation in RQ1, and further suggests that average fixation duration may be dependent on each participant’s reading and comprehension speed.
On the other hand, the conciseness of summaries is inversely correlated to the code context coverage by the programmer. From Table V, we observe that the summaries of high conciseness correspond to a significantly ($p<0.015$) lower number of fixations, lines visited, and methods visited. Figure 3 shows this delta more clearly. This suggests that participants included unnecessary information at a higher rate when they covered an exceptionally larger part of the code context. As shown in Figure 3, we did not observe a significant difference in average fixation duration or regression rate between the two groups.
Our interpretation of these observations is that there are diminishing returns in terms of increasing code context coverage and improvement in the quality of summaries. We found that gaze metrics for the summaries that achieved a perfect score for all four qualities ($Overall=5$), tend to fall between the thresholds for high completeness and conciseness. For example, in terms of lines visited, the summaries with a perfect score have a mean score of 41 lines, which is between the scores of 43.25 and 38.49 reported in Table V.
In contrast, summaries marked for low completeness have a lower mean score of 29.88, while summaries marked for low conciseness have a high mean score of 53.62, indicating that reading more context does not always result in a better summary. Rodeghero et al. [22] and Abid et al. [23] suggested that function signatures or call terms (such as method name and parameter) are key components of code in the summarization process, indicating some lines of code that contain more information than others, aligning with our finding that code context is useful for complete summaries, but concision decreases with lines read.
@@ 5.3 RQ3: Attention to Context
We found that of the types of context, participants spent a vast majority of their time looking at contextual methods and variable declarations in the same class as the target method. Surprisingly, we found that participants spent a comparatively small amount of time looking at methods in the call graph, even though we designed the second and third sentences in the summary to contain information from the callees and callers respectively. In Figure 2 we report four graphs, each showing averages for total time spent fixating on tokens inside each type of context, normalized by total time spent looking at context for the task.
In Figure 2 (a), we report the attention grouped by the project type. We found see similar trends for four of the five project, namely Scrimage, Openaudible, MLTK, and FreeCol. For these projects, participants spent an average of 40-65% of their time looking at methods in the same class as the target method. Surprisingly, for the Mallet project, on an average participants spent more time looking at the variable, declarations, and other data inside the class than methods and calls. One possible explanation for this outlier we investigated was that Mallet relies more on class variables than the other projects. However, we did not find significant differences between the number of class declarations and variables for Mallet, compared to other projects.
In Figure 2 (b), we analyze the difference in attention grouped by session, i.e., the order in which the project was processed by a participant. We found that on average as participants processed more sessions, they relied more on information inside the class, both methods and class declarations. This is expected because as participants repeat the task, they form a consistent pattern of navigating the context. This supports the earlier observation that overall, the participants found the class of the target method to be most important part of the context.
@@
In Figure 2 (d), we analyze the difference in attention of each participant. We found that most of the participants relied on class and instance methods and the class declaration for contextual information. We found that only participant 5 and 6 are outliers, and spent a significant amount of time fixating on context in other files, and these are areas that do not fall under the methods in the call hierarchy. This suggests that observations from (a), (b), and (c) are consistent with a majority of the participants and not skewed heavily by an extreme minority.
@@ 6 Example
mallet.types.Multinomial.java.ensureCapacity [line 290]
     protected void ensureCapacity (int index)      {    if (index > size)  size = index;  if (counts.length <= index) {  @Var  int newLength = ((counts.length < minCapacity)           ? minCapacity           : counts.length);  while (newLength <= index)  newLength *= 2;  double[] newCounts = new double[newLength];  System.arraycopy (counts, 0, newCounts, 0,                          counts.length);  this.counts = newCounts;  }  }      Participant 3: Accuracy: 5, Completeness: 5, Conciseness: 5, Clarity: 5
@@
Participant 4: Accuracy: 5, Completeness: 4, Conciseness: 2, Clarity: 3
     This method ensures that a double array counts has the  correct capacity for further use in estimation operations.  The method receives an index and checks if the double array  needs to be resized because the index exceeds the current  size. If it does exceed, a new array is created with the  new size and an array copy operation takes place before  setting the original counts array to the temporary one.  This utility function is useful for determining when an  array needs to be resized with further capacity. By  checking capacity before resizing, the program can ensure  memory is not being wasted
Participant 10: Accuracy: 5, Completeness: 5, Conciseness: 4, Clarity: 5
@@
Figure 4: Example with source code of a method in project mallet, accompanied by summaries written by participant 3, 4, and 10.
In Figure 4 we show an example containing target method and three corresponding summaries written by participant 3, 4, and 10. These may explain some of our observations from Section 5. Notice, the summaries written by Participant 3 and 10 are similar, but the summary written by 10 is slightly clearer. On the other hand, the summaries written by Participant 4 for the same method are lengthy and hard to comprehend and received one of the lowest overall scores by our graders (Overall = 3.5).
We found that the key to writing the last sentence for both participant 3 and participant 10’s summaries is to track the class variable minCapacity. Instances like this might explain why mallet shows a higher attention on class declarations than methods as observed in Section 5.3.
For this task, 90% of the participants first looked at the ternary operator and arraycopy elements, before looking through the call graph and class methods. We provide a representative example from participant 3: The participant started with minCapacity, before following the call graph to incremen, which is also a class method. After tracking through several overloaded methods, they looked back at the lower half of minCapacity before looking back to the instructions, where they write their summary. After looking back and forth from the target method and the call graph twice more, they move to the next task.
We also notice that participant 4’s summary is an example of summaries where excessive code coverage leads to verbose and unclear summary. Note this is just one example we found that fits the observations from Section 5.2, we do not claim that this example is representative of a trend or make any recommendations based on this example.
@@ 7 Threats to Validity
Our selection of methods is a threat to validity because, like all studies of this nature, it is not feasible to evaluate all methods and our selection of methods ultimately relies on our judgment (see Section 3.6). A different set of methods could lead to different results. The methods chosen could be too simple and not rely on context to understand them. To mitigate this threat, we chose methods that were called multiple times or had roles evident from the context. We also asked the participants in their instructions to explain the method in context of the whole program to encourage them to explore enough of the context to be accurate.
Another threat lies with our pool of participants. Our participants are graduate students, and their experiences may not translate directly to those of active industry professionals. To mitigate this threat, we screened participants with previous (and recent) Java development experience. However, industry professionals may have specific mental models for their industry and may utilize visual attention strategies that our participants do not represent. Additionally, the participants in our study had a range of experience between two and five years, meaning that it is possible that our results would not generalize to people with experience levels outside this range. Another threat to validity is our sample of participants. Even though we had 60 hours of total programmer time, these hours were from ten programmers. A different set of programmers could in theory lead to a different set of conclusions. Another threat with our pool of participants is that half of them wear corrective glasses. In specific circumstances, corrective glasses can lead to false fixation readings. To try to mitigate this threat, we recommended that participants complete the task without glasses or use contact lenses if they are able.
@@
Another threat lies with the study conditions. The room in which we conducted the study has a window. As the study was conducted over several months, the lighting conditions may have changed between sessions. Newer generation eye tracker claim to work in all lighting conditions but can cause inconsistencies in rare situations. These external threats exist for all eye tracking studies.